# EmoSphere++: Emotion-Controllable Zero-Shot Text-to-Speech via Emotion-Adaptive Spherical Vector

### v1

@@ IV-A Experimental Setup
IV-B Implementation Details
For the acoustic model, we followed the Matcha-TTS [23] configuration, utilizing the text encoder and duration predictor in the encoder, along with a 1D U-Net based CFM decoder. The CFM decoder consists of two downsampling blocks, followed by two midblock and two upsampling blocks, each containing a Transformer layer with a hidden dimensionality of 256, an attention module with dimensionality of 64, and “snakebeta” activations. Following [62], the text encoder was modified by incorporating relative position representations and adding a residual connection to the encoder pre-net. Based on [63], the duration predictor consists of two convolutional layers with rectified linear unit activation, followed by layer normalization, dropout, and a projection layer. Regarding the SER444huggingface.co/wav2vec2-large-robust-12-ft-emotion-msp-dim, we adopt a system proposed in [48], which predicts VAD using wav2vec 2.0 [64] and a linear predictor. We trained the generator using random segments of 32 frames from the Mel-spectrogram, with a batch size of 32 and a total of 11M training steps. The AdamW optimizer was used, with a $1\times 10^{-4}$ learning rate. In the inference stage, the guidance level $\gamma$ was set to 100. We trained the vocoder using the official BigVGAN [^5] https://github.com/NVIDIA/BigVGAN[61] implementation, incorporating LibriTTS [65], voice cloning toolkit [^6] https://datashare.ed.ac.uk/handle/10283/2651, and ESD datasets. All comparison models were trained using a single NVIDIA RTX A6000 GPU.
IV-C Evaluation
IV-C1 Subjective Metrics
We conducted a MOS evaluation for naturalness (nMOS), speaker similarity (sMOS), and emotion similarity (eMOS) using a nine-point scale ranging from 1 to 5, with increments of 0.5 units. The results are presented with a confidence interval of 95%. We randomly selected 10 utterances from the evaluation set.
We also conducted a preference test to evaluate the modeling of emotion expressiveness. To demonstrate the success of our modeling, we synthesized speech with three different levels of emotion intensity (weak, medium, and strong) from pairs of speech that share the same emotion and style. Evaluators were presented with two different sentences with the same emotion and style, each with varying intensities, and tasked with selecting the one exhibiting a stronger emotion. We uniformly referred to scores 0.1 as weak, 0.5 as medium, and 0.9 as strong and randomly choose 10 utterances from the evaluation set.
IV-C2 Objective Metrics
[…]
@@ IV-C2 Objective Metrics
Figure 5: Pitch tendency track according to intensity for different emotions. Pitch values were calculated by averaging the synthesized speech for each intensity across all test sentences.
@@
Moreover, we propose the SVAS to evaluate the emotion of synthesized speech. The SVAS is obtained by computing the cosine similarity of the angle of the emotion spherical vector between the synthesized audio and reference audio with the target emotion. We used a SER model101010huggingface.co/wav2vec2-large-robust-12-ft-emotion-msp-dim [48] to predict the VAD values and applied a Cartesian-to-spherical transformation through fixed neutral center coordinates to extract the angle of the emotion spherical vector.
IV-D Comparison Models
  • GT and BigVGAN [61]: The ground truth audio and waveforms are generated from the ground truth Mel-spectrogram using vocoder.
  • Mellotron [44]: This auto-regressive multi-speaker TTS model for direct style control by conditioning on rhythm and pitch utilizes conditioning GSTs.
  • Mixedemotion [10]: It is a relative attribute ranking-based model pre-computes intensity values for mixed emotion synthesis and allows manual control at run time.
  • YourTTS [46]: It is a zero-shot multi-speaker TTS with the pre-trained speaker encoder. Unlike other comparison models, it does not use a pre-trained vocoder through end-to-end training.
  • GenerSpeech [47]: It is a high-fidelity zero-shot style transfer method for out-of-domain TTS. We used the emotion2vec model instead of the fine-tuned wav2vec 2.0 model to capture global style for a fair comparison.
  • iEmoTTS [45]: This non-autoregressive TTS model for a cross-speaker emotion transfer system has been developed based on timbre-prosody disentanglement.
  • EmoSphere++: This is the proposed CFM-based emotion-controllable zero-shot TTS with EASV.
@@ IV-D Comparison Models
V Results
V-A Analysis of Emotional Style and Intensity Modeling
As previously mentioned, we characterized the derivative states of emotions through emotion style and intensity. Table I lists the results of the analysis of the novel approach to modeling emotion style and intensity using EASV. In this study, we divided the emotion space into eight regions (“I”$\sim$“VIII”) based on the VAD axes and then shifted the style along spherical spaces. Here, valence represents the positivity or negativity of an emotion, arousal indicates the intensity of the emotion provoked by a stimulus, and dominance denotes the level of control exerted by the stimulus. We aimed to demonstrate the modeling of diverse emotions by analyzing average prosodic variation rates and prosodic variation with intensity based on the characterization of the VAD axis.
TABLE V: Ablation study on disentangling modules for non-parallel emotion transfer on the seen and unseen datasets.
Method Seen Speaker Unseen Speaker
nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) EECS ($\uparrow$) nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) EECS ($\uparrow$)
GT 4.06$\pm$0.06 4.15$\pm$0.05 4.11$\pm$0.05 0.7358 0.9264 0.8311 95.53 0.9487 4.02$\pm$0.05 4.04$\pm$0.05 4.08$\pm$0.06 0.7725 0.9282 0.8504 100.00 0.9944
BigVGAN 3.95$\pm$0.05 4.01$\pm$0.06 3.98$\pm$0.06 0.7271 0.9231 0.8251 94.25 0.9389 3.93$\pm$0.06 4.01$\pm$0.05 4.04$\pm$0.06 0.7603 0.9275 0.8439 99.32 0.9903
w/o Disentangling Module 3.88$\pm$0.06 3.97$\pm$0.06 3.82$\pm$0.07 0.7289 0.9052 0.8171 92.86 0.9218 3.88$\pm$0.06 3.93$\pm$0.06 3.82$\pm$0.07 0.6430 0.8715 0.7573 88.56 0.8802
w/ Orthogonality Loss [15] 3.86$\pm$0.06 3.95$\pm$0.06 3.83$\pm$0.06 0.7287 0.9032 0.8160 92.26 0.9191 3.92$\pm$0.06 3.93$\pm$0.06 3.81$\pm$0.07 0.6454 0.8644 0.7549 91.23 0.9101
w/ Vector Quantization [47] 3.91$\pm$0.06 3.96$\pm$0.06 3.80$\pm$0.07 0.7258 0.9027 0.8143 92.34 0.9235 3.85$\pm$0.06 3.96$\pm$0.06 3.82$\pm$0.08 0.6170 0.8354 0.4177 92.60 0.9320
EmoSphere ++ (Proposed) 3.92$\pm$0.06 3.97$\pm$0.06 3.86$\pm$0.06 0.7314 0.9047 0.8181 93.53 0.9270 3.92$\pm$0.06 3.99$\pm$0.05 3.85$\pm$0.07 0.6543 0.8640 0.7592 94.61 0.9385
V-A1 Prosodic Variation with Intensity Based on the Valence-Arousal-Dominance (VAD) Axis
In this section, we describe the analysis of how prosodic features, such as pitch, vary with emotion intensity based on the VAD axis. We represent the region divided into thirds by intensity with Q1, Q2, and Q3, using 0.33 and 0.66 as thresholds. By observing the number of instances modeled, we observe that the spherical vectors representing each emotion concentrate on the unique styles frequently associated with each emotion. This pattern reflects how people tend to express certain emotion styles more regularly in reality. For instance, in the case of surprise, a higher pitch is more commonly expressed than a lower pitch.
We observe that the prosodic pattern based on emotional style and intensity reflects the characteristics of the VAD axis. Positive valence is associated with higher pitch and energy as the emotional intensity increases, whereas the duration tends to shorten. Similarly, positive arousal leads to increased patterns of pitch and energy change but typically decreases patterns of change in duration. By contrast, with positive dominance, the pitch and energy variation ranges become narrower, whereas the duration tends to be broader. For example, in the happy emotion, we can observe a shift in the range between Q1 and Q3 as dominance changes, leading to distinct variations in “I” and “V” styles. Similarly, the “VII” style in the sad emotion reflects negative arousal, with prosodic variation patterns changing with intensity. Compared to the “VIII” style, the difference in valence results in noticeable differences in the average prosodic patterns.
V-A2 Average Prosodic Variation Rates
To demonstrate the diverse modeling of emotions, we used a metric that quantifies the overall emotion averages for prosody, energy, and duration and the average prosodic variation rates for the corresponding emotion styles. This metric $p_{c}$ is calculated using the formula $p_{s}-p_{a}/p_{a}$, where $p_{s}$ represents the average prosodic feature value for a specific emotion style and $p_{a}$ denotes the overall average prosodic feature value for that emotion. We observed different combinations of prosodic variation rates $p_{c}$ of pitch, energy, and duration for each emotional style. For example, we can observe that the cell hue combinations in the average prosodic variation rates vary across different emotions, highlighting the diversity in emotional expression. In summary, the results are consistent with our assumptions about speech emotion analysis, underscoring the efficacy of our modeling in accounting for emotional style diversity.
V-B Model Performance
We conducted experiments including seen and unseen scenarios to evaluate EmoSphere++ and baseline models for non-parallel style transfer. We split our experiments into two categories: 1) seen non-parallel style transfer and 2) unseen non-parallel style transfer.
@@ V-C1 Comparison With Control Methods
As a comparative study, we implemented two speaker and emotion attribute control methods (i.e., Matcha-TTS w/ Scaling Factor, Matcha-TTS w/ Relative Attributes), as described in Section IV-D. We evaluated the performance of these two methods in terms of performance and control. To demonstrate the ability of our model to control intensity, we synthesized speech with three different levels of emotion intensity (weak, medium, and strong). In the relative attribute model and EmoSphere++, we uniformly refer to values of 0.1 as weak, 0.5 as medium, and 0.9 as strong. The scaling factor cannot assign intensity values; therefore, we set the scalar factor to 1, 2, and 3 to represent weak, medium, and strong emotion intensities, as in the original setting.
@@
Control. We calculated the average pitch of the synthesized speech across all test sentences based on the intensity of each emotion. Fig. 5 shows the following: 1) The Matcha-TTS w/ Relative Attributes reflects the adjustment of emotional properties, but it often shows a limited adjustable range and tends to be reduced to a more uniform style. This outcome suggests subtle emotional nuances cannot be easily captured using only emotion labels. 2) The Matcha-TTS w/ Scaling Factor exhibits the most variation, but it often becomes unstable when adjusted on labels such as sad. Therefore, determining an appropriate scaling factor is difficult, and adjustments may lead to instability in audio quality. Conversely, the pitch tendency plot of EmoSphere++ reflects the intensity based on emotion, indicating that the proposed model synthesizes speech according to the given intensity scale.
@@ V-C2 Intensity Control in the Zero-Shot Scenario
Figure 9: Pitch tendency track according to intensity for each emotion. Pitch values were calculated by averaging the synthesized speech for each intensity across all test sentences.
V-D Emotion Style Shift
V-D1 Visual Comparisons
We visualized the prosodic attributes of pitch related to the emotion intensity to gain an intuitive understanding of emotion style. To illustrate the variation patterns in emotion intensity with shifted emotion style, we visualized the pitch track changes for a sample, both from seen and unseen speakers. As analyzed in Section V-A, the prosodic pattern based on emotional intensity reflects the characteristics of the AVD axis. As shown in Fig. 7, we visualized the pitch contour of sad utterances with the most varied styles. For example, style vectors with positive V axes have higher average pitch values and reduced duration; positive A axes have a pitch that tends to increase the changing patterns. By contrast, duration decreases, and positive D axes have a narrow range in changing patterns with a broader duration. These results indicate that the proposed EASV is meaningfully characterized and enables emotion style and intensity controllable speech synthesis in both seen and unseen.
V-D2 Comparison of Emotional Consistency
V-E Ablation Study
V-E1 Impact of Disentangling Module
We conducted ablation studies to evaluate the impact of the disentangling module used in EmoSphere++ and present the findings in Table V. We used comparative subjective metrics (nMOS, sMOS, and sMOS) and objective metrics (SECS, ECA, and EECS) to assess the expressiveness and quality of the generated speech. The increase in quality and expressiveness demonstrates the importance of disentangling the module between emotion and speaker embeddings. We also determined that our proposed normalized orthogonality loss significantly increases the expressiveness of the unseen case, indicating that it has an important role in providing strong generalization performance for zero-shot scenarios.

### v2

@@ IV-D Comparison Models
V Results
V-A Analysis of Emotion-Adaptive Spherical Vectors for Emotional Style and Intensity
As previously mentioned, we characterized the derivative states of emotions through emotion style and intensity using the emotion-adaptive spherical vector (EASV). To evaluate the effectiveness of modeling based on emotion style and intensity, we conducted an analysis using the ESD dataset along with large-scale datasets, including MSP-Podcast corpus [64] and IEMOCAP datasets [65]. In this study, we divided the emotion space into eight regions (“I”$\sim$“VIII”) based on the VAD axes and then shifted the style along spherical spaces, using only the five categorical emotion labels from the ESD dataset for consistency. As shown in Table I, we illustrate the prosodic variation and distribution of large-scale emotional speech datasets based on emotion style and intensity modeling. The analysis details are described in the following subsections.
V-A1 Analysis of Emotion Dataset Distribution by Style and Intensity
To demonstrate the diverse modeling of emotions, we analyzed the distribution of emotion styles and intensities using EASV modeling. We observe that each emotion exhibits a range of styles and intensities, reflecting the natural variability in emotional expression. By analyzing the number of modeled instances, we find that the spherical vectors representing each emotion tend to cluster around specific styles that are frequently associated with that emotion. This suggests that certain emotion styles are more commonly expressed in real-world speech. For example, in the case of surprise, a higher pitch is more frequently observed than a lower pitch. However, for styles with fewer data instances, the distribution showed less distinct patterns, likely due to the limited availability of data. In summary, the results support the assumption that speech emotion analysis should account for diverse styles based on primary emotional states.
@@ V-A1 Analysis of Emotion Dataset Distribution by Style and Intensity
Figure 5: Pitch tendency track according to intensity for different emotions. Pitch values were calculated by averaging the synthesized speech for each intensity across all test sentences. Since the intensity of ground truth (GT) speech cannot be adjusted, the GT line represents the pitch tendency based on the emotion-adaptive spherical vector intensity labels across all test sentences, serving as a reference guideline.
@@
V-A2 Prosodic Variation with Intensity Based on the Valence-Arousal-Dominance (VAD) Axis
In this section, we describe the analysis of how prosodic variation varies with emotion intensity based on the VAD axis. We represent the region divided into thirds by intensity with Q1, Q2, and Q3, using 0.33 and 0.66 as thresholds. In psychology [1, 36], valence represents the positivity or negativity of an emotion, arousal indicates the intensity of the emotion provoked by a stimulus, and dominance denotes the level of control exerted by the stimulus. Building on this, we show how prosodic variations reflect diverse emotions through emotion style and intensity along the VAD axis.
Existing studies [11] on emotion control have demonstrated the effectiveness of analyzing prosodic features such as pitch, energy, and duration. Expanding on this approach, we conducted a prosodic analysis and found that: 1) positive valence is associated with higher prosodic feature value as the emotional intensity increases. 2) positive arousal leads to increased patterns of prosodic feature value change. 3) positive dominance results in a narrower prosodic variation range. To further analyze and clearly illustrate this prosodic variation, we introduced three key elements in our analysis. First, to compare differences based on valence, we calculated the average prosodic feature values AVG. for each emotion style. Second, to observe increases and decreases related to arousal, we highlighted the highest and lowest prosodic feature values within each intensity level using green and red colors, respectively. Lastly, to examine the variation ranges influenced by dominance, we introduced $Qc$, which represents the absolute difference between the highest and lowest prosodic feature values. These measures provide a more comprehensive understanding of how prosodic features vary with emotional style and intensity.
To validate these findings, we examined specific cases where emotion styles differ along a single VAD dimension while remaining constant in the others. First, when comparing styles “III” and “IV”, where only valence differs, we observe that positive valence is associated with higher average prosodic feature values AVG.. Specifically, across all emotions, converting to positive valence results in an average increase of 13.45 in pitch, 1.4 in energy, and 0.25 in duration. Similarly, when comparing styles “I” and “V”, where only dominance differs, we find that $Qc$ is smaller in the positive dominance condition. Specifically, converting to positive dominance results in an average decrease of 6.78 in pitch, 0.35 in energy, and 0.02 in duration across all emotions. Lastly, the positive arousal style of “I” exhibits increasing prosodic patterns, whereas the negative arousal style of “IV” shows decreasing patterns. These results show that prosodic variations align with the VAD characterization in most cases, supporting the effectiveness of EASV in modeling emotion style and intensity.
V-B Model Performance
We conducted experiments including seen and unseen speaker scenarios to evaluate EmoSphere++ and baseline models for non-parallel style transfer. We split our experiments into two categories: 1) seen non-parallel style transfer and 2) unseen non-parallel style transfer.
@@ V-C1 Comparison With Control Methods
As a comparative study, we implemented two speaker and emotion attribute control methods (i.e., Matcha-TTS w/ Scaling Factor, Matcha-TTS w/ Relative Attributes), as described in Section IV-D. We evaluated the performance of these two methods in terms of performance and control. To demonstrate the ability of our model to control intensity, we synthesized speech with three different levels of emotion intensity (weak, medium, and strong). In the relative attribute model and EmoSphere++, we uniformly refer to values of 0.1 as weak, 0.5 as medium, and 0.9 as strong. The scaling factor cannot assign intensity values; therefore, we set the scalar factor at 1, 2, and 3 to represent weak, medium, and strong emotion intensities, as in the original setting. The GT line reflects the pitch tendency of the original speech, based on inherent EASV intensity labels across all test sentences, serving as a reference guideline.
@@
Control. We calculated the average pitch of the synthesized speech across all test sentences based on the intensity of each emotion. Fig. 5 shows the following: 1) The Matcha-TTS w/ Relative Attributes reflects the adjustment of emotional properties, but also often shows a limited adjustable range and tends to be reduced to a more uniform style. This outcome suggests subtle emotional nuances cannot be easily captured using only emotion labels. 2) The Matcha-TTS w/ Scaling Factor exhibits the most variation, but often becomes unstable when adjusted on labels such as sad. Therefore, determining an appropriate scaling factor is difficult, and adjustments may lead to instability in audio quality. Conversely, the pitch tendency plot of EmoSphere++ closely follows the GT line, reflecting intensity variations based on emotion. This result indicates that the proposed model synthesizes speech according to the given intensity scale while effectively capturing variations that align more closely with natural emotional speech patterns.
TABLE V: Ablation study on disentangling method and joint attribute style encoder for non-parallel emotion transfer on the seen and unseen datasets. The nMOS, sMOS, and eMOS scores are presented with 95% confidence intervals.
Method Seen Speaker Unseen Speaker
nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) EECS ($\uparrow$) nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) EECS ($\uparrow$)
GT 4.23$\pm$0.04 4.19$\pm$0.03 4.08$\pm$0.03 0.8311 95.53 0.9487 4.16$\pm$0.05 4.11$\pm$0.05 4.07$\pm$0.03 0.8504 100.00 0.9944
BigVGAN 4.22$\pm$0.04 4.13$\pm$0.04 4.05$\pm$0.04 0.8251 94.25 0.9389 4.12$\pm$0.06 4.05$\pm$0.04 4.02$\pm$0.04 0.8439 99.32 0.9903
w/o Disentangling Method 3.83$\pm$0.05 3.78$\pm$0.05 3.75$\pm$0.04 0.8171 92.86 0.9218 3.70$\pm$0.08 3.83$\pm$0.05 3.70$\pm$0.04 0.7573 88.56 0.8802
w/o Dimensional Emotion Encoder 3.79$\pm$0.05 3.75$\pm$0.04 3.74$\pm$0.05 0.8165 92.18 0.9272 3.73$\pm$0.08 3.81$\pm$0.04 3.72$\pm$0.05 0.7574 88.27 0.8900
w/o Global Emotion Encoder 3.76$\pm$0.05 3.77$\pm$0.04 3.51$\pm$0.05 0.8178 77.68 0.8171 3.64$\pm$0.08 3.79$\pm$0.05 3.60$\pm$0.04 0.7568 59.66 0.6712
w/ Gradient Reversal Layer [22] 3.82$\pm$0.05 3.76$\pm$0.04 3.77$\pm$0.04 0.8169 92.76 0.9283 3.75$\pm$0.07 3.81$\pm$0.05 3.71$\pm$0.04 0.7566 89.58 0.8947
w/ Vector Quantization [47] 3.78$\pm$0.05 3.74$\pm$0.03 3.76$\pm$0.05 0.8143 92.34 0.9235 3.76$\pm$0.08 3.67$\pm$0.05 3.74$\pm$0.04 0.7262 92.60 0.9320
w/ Orthogonality Loss [15] 3.79$\pm$0.06 3.76$\pm$0.04 3.76$\pm$0.04 0.8160 92.26 0.9191 3.74$\pm$0.08 3.74$\pm$0.05 3.73$\pm$0.05 0.7549 91.23 0.9101
EmoSphere ++ (Proposed) 3.87$\pm$0.05 3.77$\pm$0.04 3.78$\pm$0.05 0.8181 93.53 0.9270 3.76$\pm$0.07 3.83$\pm$0.05 3.75$\pm$0.04 0.7592 94.61 0.9385
@@ V-C2 Intensity Control in the Zero-Shot Scenario
V-D Emotion Style Shift
V-D1 Visual Comparisons
We visualized the prosodic attributes of pitch related to the emotion intensity to gain an intuitive understanding of emotion style. To illustrate the variation patterns in emotion intensity with shifted emotion style, we visualized the pitch track changes for a sample, both from seen and unseen speakers. As analyzed in Section V-A, the prosodic pattern based on emotional intensity reflects the characteristics of the VAD axis. As shown in Fig. 7, we visualized the pitch contour of sad utterances with the most varied styles. For example, style vectors with positive V axes have higher average pitch values and reduced duration; positive A axes have a pitch that tends to increase the changing patterns. By contrast, duration decreases, and positive D axes have a narrow range in changing patterns with a broader duration. These results indicate that the proposed EASV is meaningfully characterized and enables emotion style and intensity controllable speech synthesis in both seen and unseen.
V-D2 Comparison of Emotional Consistency
V-E Ablation Study
V-E1 Impact of Disentangling Method and Joint Attribute Style Encoder
As shown in Table V, we conducted ablation studies to evaluate the impact of the disentangling method and joint attribute style encoder. To ensure a fair comparison of the proposed joint attribute style encoder modules and the existing disentangling approaches, we trained the models using the same dataset and configurations as EmoSphere++. We replaced the existing emotion disentangling method with three competing disentangling approaches: 1) w/ Gradient Reversal Layer [22], which employs adversarial speaker training using the GRL [50] applied after the fully connected layer that follows the global encoder; 2) w/ Vector Quantization [47], which implements a bottleneck layer via a modified VQ layer [51] applied after the fully connected layer that follows the global encoder; and 3) w/ Orthogonality Loss [15], which constrains the emotion embedding and speaker embedding using an orthogonal loss [52] without normalizing all sample pairs. To further evaluate the effectiveness of the joint attribute style encoder, we conducted additional ablation studies by removing individual encoder components: w/o Global Emotion Encoder, where the global emotion embedding is excluded and w/o Dimensional Emotion Encoder, where the dimensional-driven emotion embedding is removed. Additionally, to analyze the role of the disentangling method, we trained a variant of the proposed model without the disentangling method, referred to as w/o Disentangling Method. In the w/o Global Emotion Encoder setting, the disentangling method was modified to use the dimensional-driven emotion embedding in place of the global emotion embedding.
@@ V-E1 Impact of Disentangling Method and Joint Attribute Style Encoder
Figure 9: Pitch tendency track according to intensity for each emotion. Pitch values were calculated by averaging the synthesized speech for each intensity across all test sentences. Since the intensity of ground truth (GT) speech cannot be adjusted, the GT line represents the pitch tendency based on the emotion-adaptive spherical vector intensity labels across all test sentences, serving as a reference guideline.
We used comparative subjective metrics (nMOS, sMOS, and sMOS) and objective metrics (SECS, ECA, and EECS) to assess the expressiveness and quality of the generated speech. The experimental results show that improvements in both quality and expressiveness highlight the effectiveness of disentangling emotion and speaker embeddings, ensuring clearer emotional transfer without compromising speaker identity. Furthermore, our findings demonstrate that integrating global and dimensional-driven features in the joint attribute style encoder enables the model to capture both broad and fine-grained characteristics, further enhancing expressive. Specifically, the proposed normalized orthogonality loss is crucial for preserving emotional expressiveness in unseen cases, reinforcing its importance for achieving strong generalization performance in zero-shot scenarios.